OpenStudy (anonymous):

solve each system by elimination Please help and solve y 36+7x-8y=0 -6x+6y=18

OpenStudy (across):

We have the following two equations: \[7x-8y+36=0\]\[-6x+6y=18\]We want to solve for y. Therefore, we multiply the first equation by 6 and the second equation by 7: \[42x-48y+216=0\]\[-42x+42y=126\]If we add these two equations, the x's cancel out and we're left with: \[-6y=342\]\[y=-57\]

OpenStudy (bme89):

Adding the two equations should give -6y=-90, giving y=15

OpenStudy (across):

... and make sure you move the constant to the other side ;p

OpenStudy (anonymous):

wait whats x

OpenStudy (bme89):

just plug y=15 into one of the equations and solve for x